OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±7).

OpenStudy (mertsj):

Is this hyperbola horizontal or vertical?

OpenStudy (anonymous):

@Mertsj vertical

OpenStudy (mertsj):

Good. That means the y term will be positive. Do you agree?

OpenStudy (anonymous):

@Mertsj yes

OpenStudy (mertsj):

And we know for a hyperbola that : \[c^2=a^2+b^2\]

OpenStudy (mertsj):

And we know that the distance from the center to the vertex is "a" and the distance from the center to the focus is "c"

OpenStudy (mertsj):

We know that this hyperbola has center (0,0)

OpenStudy (mertsj):

So its equation will look like this: \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\]

OpenStudy (mertsj):

So now, you should be able to write its equation.

OpenStudy (anonymous):

@Mertsj my only problem is I don't know what #'s I should plug in

OpenStudy (mertsj):

What is the center of this hyperbola?

OpenStudy (anonymous):

@Mertsj (0,0)?

OpenStudy (mertsj):

Yes. Now you need to think about the vertex. What is the vertex?

OpenStudy (anonymous):

@Mertsj I don't know

OpenStudy (mertsj):

Allow me to suggest that you read the problem.

OpenStudy (anonymous):

@Mertsj the vertex is (0,2)

OpenStudy (mertsj):

How far is it from the center to the vertex?

OpenStudy (anonymous):

@Mertsj 2 units

OpenStudy (mertsj):

Do you remember that I said the distance from the center to the vertex is "a"?

OpenStudy (anonymous):

@Mertsj so its 4

OpenStudy (mertsj):

a^2=4. Yes.

OpenStudy (anonymous):

if it is 4 then b = 49

OpenStudy (mertsj):

No.

OpenStudy (mertsj):

What is the focal point?

OpenStudy (anonymous):

@Mertsj its 7

OpenStudy (mertsj):

According to your problem, the focal points are (0,7) and (0,-7). That tells you that the focal points are 7 units from the center and that c =7. Use the relationship between a, b, and c to find b.

OpenStudy (anonymous):

@Mertsj thanks for the enlightenment it will be 7^2=2^2+b^2 = 49=4+b : 49-4= 45

OpenStudy (mertsj):

Perfect!! Now can you write the equation? What is it?

OpenStudy (anonymous):

@Mertsj \[\frac{ y ^{2} }{ 4 }-\frac{ x ^{2} }{ 45 }=1\]

OpenStudy (mertsj):

Wonderful!! You get an A for the day!!

OpenStudy (anonymous):

@Mertsj thanks for the help

OpenStudy (mertsj):

You're welcome. (I love conic sections)